What is the power in diopters of converging lens of focal length 0.5m, 2m, 0.5cm


1
Expert's answer
2021-06-11T11:20:46-0400

"D=\frac 1F,"

"D_1=\frac{1}{0.5}=2~dpt,"

"D_2=\frac{1}{2}=0.5~dpt,"

"D_3=\frac{1}{0.05}=20~dpt."